The equation of the line through (8,6) and (2,-4) is?
Answer by checkley75(3666)

Slope=(-4-6)/(2-8)=-10/-6=5/3
Now substitute one set of x,y values in the line equation [Y=mX+b].
WHERE (m)=SLOPE & (b)=Y INTERCEPT.
6=5/3*8+b
6=40/3+b
b=6-40/3
b=(18-40)/3
b=-22/3
Y=5X/3-22/3 ANSWER.